OpenStudy (anonymous):

How do I write this question, 2^1/3 in the radical form?

OpenStudy (anonymous):

\[^3\sqrt{2}\]

OpenStudy (anonymous):

If your exponent is a fraction, the denominator is the root and the numerator is the number outside of your root.

OpenStudy (anonymous):

In general the relationship between rational exponent and radical is\[x ^{\frac{m}{n}}=\sqrt[n]{x^m}\]

OpenStudy (anonymous):

as long as n != 0

OpenStudy (anonymous):

what if n < 0?

OpenStudy (anonymous):

yes you are correct!

OpenStudy (anonymous):

Oh how about if the exponent is negative? Such as: 7^-1/2?

OpenStudy (anonymous):

negative exponent means reciprocal regardless if the exponent is integer, rational or real.

OpenStudy (anonymous):

so you mean, the negative exponent would be 2/1?

OpenStudy (anonymous):

\[7^{-1/2}=\frac{1}{7^{1/2}}\]

OpenStudy (anonymous):

reciprocal with respect to the base, not the exponent

OpenStudy (anonymous):

ohhh, thank you!

OpenStudy (anonymous):

You do know that in radical form, a radical in the denominator is not considered simplified. If you are in a class below calculus, you will have to "rationalize" the denominator.

OpenStudy (anonymous):

\[=\frac{1}{\sqrt{7}}*\frac{\sqrt{7}}{\sqrt{7}}=\frac{\sqrt{7}}{7}\]